OpenStudy (anonymous):

I will award you! :) Solve the system of equations using elimination. Write your answer as (x, y). 2x-3y=7 x+2y=14 A. (8,3) B. (-3,-4 1/3) C. (-28,21) D. (1,-1 2/3)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

A is the answer

OpenStudy (anonymous):

1. 2x-3y=7 x+2y=14 2. 2(8)-3(3)=7 (8)+2(3)=14 3. 16-9=7 8+6=14 4. A(8,3) the answer is A. and that's my proof